# sponchpop II: Population Synthesis to Investigate Volatile Sulfur as a Fingerprint of Gas Giant Formation Histories

### v1

@@ Abstract
Planet population synthesis is an integral tool for linking exoplanets to their formation environments. Most planet population synthesis studies have focused on the carbon-to-oxygen ratio (C/O) in gas or solids, yet more insight into planet formation may be afforded by considering a wider suite of elements. Sulfur is one such key element. It has been assumed to be entirely refractory in population synthesis models, restricting it to being a tracer of accreted rocky solids. However, sulfur also has a volatile reservoir dominant at the onset of star and planet formation. We investigate sulfur’s wider potential as a formation history tracer by implementing the first multi-phase treatment of S in a planet population synthesis model. We present the planet formation module of sponchpop and its first predicted planet growth tracks and populations. We explore the diversity of planet compositions in terms of their sulfur budget, including both refractory and volatile components, and apply a novel gas-grain conversion of sulfur to study how formation trajectories of giant planets relate to final core and envelope compositions. We show that planets inherit a wide range of core and envelope sulfur content related to accretion history while considering late-stage planetesimal infall, providing a new diagnostic tool for planet formation. The diverse sulfur content of planet cores suggests some rocky planets may be born sulfur-poor, with implications for their geochemistry and habitability. Enhanced sulfur abundances in gas-giant atmospheres can be attributed to formation beyond the $\mathrm{H_{2}S}$ iceline, such as the giants in our Solar System.
@@
pubyear: 2025pagerange: sponchpop II: Population Synthesis to Investigate Volatile Sulfur as a Fingerprint of Gas Giant Formation HistoriesA
@@ 1 Introduction
As of writing, just under 6000 exoplanets have been discovered, the diversity of which far exceeds that of the planets of our own Solar System. The properties of these exoplanetary bodies, their masses, semi-major axes, chemical compositions, and the systems they are part of all contribute to the understanding of their formation processes. While these formation processes may be common, diversity arises from the differences in starting conditions, such as ‘birth’ locations, host star properties, initial disk conditions and encounters with other planets in the same systems. Planet population synthesis modelling can bridge the gap between observations of protoplanetary disks and the fully formed planetary systems seen today (Mordasini et al., 2009; Benz et al., 2014).
Combining the data obtained from the Atacama Large Millimetre/sub-millimetre Array (ALMA) and the James Webb Space Telescope (JWST), which details the chemical composition of planet forming environments and the bulk chemical abundances of the the atmospheres of gas giants, enables us to link protoplanetery disks to the planets they produce. The S-, P-, O-, N- and C-bearing species, when observed in the atmospheres of giant planets, can provide information on planets’ atmospheric processes and provide valuable insights about possible formation histories (Öberg and Wordsworth, 2019).
@@
In the ISM, sulfur exists almost entirely in atomic form, but in protoplanetary disks, it partitions between a small set of gas-phase molecules (e.g. $\mathrm{H_{2}S}$, CS, SO) and refractory condensates (such as FeS, MgS) (Semenov et al., 2018; Kama et al., 2019). The form sulfur takes depends on the local temperature and pressure conditions in a disk (Fig. 1).
[…]
Figure 1: Fractional representation of volatile and refractory reservoirs of sulfur throughout the disk evolution and planet formation processes. Values are normalised to solar. (Kama et al., 2019)
@@
Hot Jupiters represent ideal laboratories for testing sulfur chemistry models due to their extreme atmospheric conditions and observational accessibility. Their close-in orbits result in atmospheric temperatures exceeding $1000\rm K$, where sulfur-bearing species achieve chemical equilibrium and become detectable through transmission and emission observations (Hobbs et al., 2021). Critically, at these temperatures, sulfur species remain in the gas phase above the photosphere, making atmospheric abundances observed in transmission spectroscopy a reliable proxy for the complete atmospheric and envelope inventory of sulfur (Fu et al., 2024). This absence of sulfur condensation at observable pressures distinguishes hot Jupiters from cooler gas giant planets where sulfur may condense below the photosphere, obscuring bulk composition. JWST’s infrared capabilities have already begun revealing substantial sulfur in both Hot Jupiter and Neptune atmospheres that cannot be explained by refractory-only models.
@@
The detection of sulfur dioxide in WASP-39b’s atmosphere by JWST represents a breakthrough in exoplanet atmospheric chemistry (Alderson et al., 2023; Tsai et al., 2024; Powell et al., 2024). JWST has also detected hydrogen sulfide ($\mathrm{H_{2}S}$) in HD189733b’s atmosphere, providing the first direct measurement of reduced sulfur species in an exoplanet (Fu et al., 2024), and sulfur dioxide in warm Neptune WASP-107b (Dyrek et al., 2024). These discoveries demonstrate that several reservoirs of sulfur in exoplanet atmospheres are detectable, and may provide new constraints on models of planet formation and atmospheric chemistry.
@@
Initial Embryo Mass $M_{p,0}$ - $10^{-5}M_{\oplus}$
@@
‘Birth’ Semi-Major Axis $a_{0}$ $0.05-r_{out,0}$ -
Planetesimal-to-pebble ratio $f_{ptp}$ - $0.1$
Fragmentation velocity $v_{frag}$ 2.5 $ms^{-1}$ -
@@
Initial disk mass $M_{d,0}$ $0.01-0.1M_{\odot}$ $1\times 10^{-2}M_{\odot}$
Disk wind strength* $f_{w}$ 010-1 $0.8$
@@ 2.1 Disk Models & Initial Conditions
The disk module in sponchpop allows for a variety of 1D, or 1+1D protoplanetary disk models to be implemented. It includes 1) a static disk set to match that used in Öberg et al. (2011); 2) viscous and irradiated disks from Chambers (2009); and 3) a viscous, irradiated disk under the influence of an magnetohydrodynamical (MHD) disk wind from Chambers (2018). For the case of the MHD wind disk model, we use fiducial parameters consistent with a fast wind and negligible mass loss due to the winds. In both the MHD wind and viscous irradiated disk models, the boundary between the inner disk (heated through viscous motion) and the flared outer disk (heated through stellar irradiation) are divided by a transition radius that moves inwards over time as the disk mass accretion rate declines. All host stars considered in this study are have solar mass, radius and effective temperature. The initial elemental abundances for $\rm Fe$ and $\rm S$ are taken from Asplund et al. (2009).
@@ 2.2.3 Migration
In the early stages of the planet formation process when the protoplanet is not yet, or never will, reach a mass significant enough to perturb the gas disk, it undergoes type I accretion, wherein the inwards drift of a disk-embedded planet is determined by the interaction between the disk and the planet. Inward migration in this case determined by an imbalanced torque from Lindblad resonances within the planet’s orbit. The migration prescription utilised is from Tanaka et al. (2002). This formulation is used as it assumes the protoplanetary disk is locally vertically isothermal, which is the case for our disk models. Using the total net torque (Lindblad and co-rotation) for a 2D disk,
@@ 2.3 Chemistry
Figure 6: Formation tracks of planets in the viscous irradiated disk with an alpha viscosity value of $\alpha=0.7\times 10^{-3}$ and starting mass of $M_{disk,0}=0.075M_{\odot}$. All other parameters are fiducial. The solid, dashed and dotted parts of each track represent the solid accretion, envelope contraction and runaway gas accretion, respectively. The marker colour denotes the final sulfur mass fraction in the envelope (outer ring) and core (solid centre) respectively, and planets with negligible sulfur content are marked in white (values below $10^{-10}$). The blue shaded region represents the minimum and maximum pebble isolation mass. Planet markers that are empty are planets that have no/negligible sulfur in both the core and envelope. This set of planets is generated with a $p_{\rm ratio}=0$, to demonstrate spatial dependence on sulfur abundance without supplementation from refractory sulfur reservoirs.
@@ 3 Results
Figure 9: Violin plots showing the non-zero sulfur mass fractions in the cores and envelopes of giant planets (0.2 - 2 $M_{\rm J}$) across all three disk models. Each row corresponds to one of the three values considered for the ratio of planetesimals ablated in the planets’ atmospheres. Two cases of $p_{\rm ratio}=1$ are featured one where sulfur is only considered in the refractory carried by planetesimals, and one where the chemical kinetics described in Section 2.3 are included, therefore including both gas-phase and refractory sulfur reservoirs. Jupiter’s atmospheric sulfur mass fraction is marked in its appropriate mass category for reference.
@@ 4.3 Final Sulfur Abundances
In the our $p_{\rm ratio}=0$ case, the maximum sulfur abundance achieved by a gas giant ($\geq 50M_{\oplus}$) was 1.4 times the present day solar abundance. The maximums reached by the $p_{\rm ratio}=0.5$ and $p_{\rm ratio}=1$ were 5.9 and 9.5 times solar, respectively. The most similar planet to Jupiter generated in our models (with a mass of $321.47\rm M_{\oplus}$ and final semi-major axis of $5.16\rm AU$) produced a final envelope sulfur abundance of 1.37 times the solar sulfur abundance. In order to achieve Jupiter’s sulfur abundance of $4.45\times 10^{-5}$, the planet would need an additional $0.0165M_{\oplus}$ of sulfur in its atmosphere, whether through bombardment of sulfur-rich planetesimals, or significant core-envelope mixing. However, Jupiter’s enhanced sulfur abundance is able to be met by comparably sized cold gas giant planets in the case of $p_{\rm ratio}=0.5$ if accreting sulfur-rich planetesimals beyond the $\mathrm{H_{2}S}$ iceline. This only true of planets that have migrated into regions of the disk that have not been depleted of sulfur rich planetesimals during the core accretion stage. This is dependant on the alpha viscosity of the disk, which heavily influences the migration speed of the planet, particularly during type II migration.
@@ 5 Conclusions
    Sulfur content in planet cores and rocky planets is sensitive to a coupled treatment of volatile and refractory sulfur and is intrinsically related to the formation and migration history. Cores born in the inner few AU ($\leq 3\rm AU$ in our fiducial model) are not supplemented efficiently by refractory sulfur even in the inner disk, as the temperature range for efficient FeS formation is conducive to a much smaller region of the disk than previously understood. Treating sulfur as only being carried in the refractory neglects the solid sulfur ‘desert’, wherein conditions for efficient iron sulfide production are not met and $\mathrm{H_{2}S}$ does not freeze out. This fundamentally changes sulfur core budget predictions for gas giant cores and terrestrial planets.
@@
    Atmospheric sulfur abundances indicate formation and migration histories for planets with substantial envelopes. Super-solar atmospheric sulfur requires formation beyond the $\mathrm{H_{2}S}$ iceline and efficient planetesimal ablation in the outer disk after the onset of runaway gas accretion. Significant sulfur depletion ($10^{-8}$) indicates either formation in volatile sulfur depleted zones due to efficient refractory S conversion, or lack of access to sulfur bearing planetesimals for late-stage infall. This formation diagnostic can compliment standard C/O analysis based approaches to isolate possible migration histories. The dependence of sulfur chemistry on the local thermal and pressure conditions of the disk midplane offers information regarding the formation histories of gas-giants. Migration histories generate core-envelope compositions that indicate possible relative timing of solid and gas accretion phases, as a coupled treatment of volatile and refractory sulfur can provide information regarding the timing of solid and gas accretion through untouched or depleted planetesimal reservoirs.

### v2

# sponchpop: Population synthesis to investigate volatile sulfur as a fingerprint of gas giant formation histories

@@ Abstract
Planet population synthesis is an integral tool for linking exoplanets to their formation environments. Most planet population synthesis studies have focused on the carbon-to-oxygen ratio (C/O) in gas or solids, yet more insight into planet formation may be afforded by considering a wider suite of elements. Sulfur is one such key element. It has been assumed to be entirely refractory in population synthesis models, restricting it to being a tracer of accreted rocky solids. However, sulfur also has a volatile reservoir dominant at the onset of star and planet formation, which is then converted into refractories. We investigate sulfur’s wider potential as a formation history tracer by implementing a gas-grain chemical conversion, the first multi-phase treatment of S in a planet population synthesis model. We also present the planet formation module of sponchpop and its first predicted planet growth tracks and populations. We apply these to explore the diversity of the planetary sulfur budget. We show that planets can inherit a wide range of core and envelope sulfur content, depending on their formation environment and accretion history including late-stage infall, demonstrating sulfur’s new potential as a diagnostic tool for planet formation. Our models predict that some rocky planets are born sulfur-poor, which may have significant implications for their geochemistry and habitability. Enhanced sulfur abundances in gas-giant atmospheres, such as in our solar system, may result not only from accretion of rocky planetesimals, but also from formation beyond the $\mathrm{H_{2}S}$ iceline.
@@
pubyear: 2026pagerange: sponchpop: Population synthesis to investigate volatile sulfur as a fingerprint of gas giant formation historiesA
@@ 1 Introduction
As of writing, just under 6000 exoplanets have been discovered, the diversity of which far exceeds that of the planets of our own Solar System. The properties of these exoplanetary bodies, their masses, semi-major axes, chemical compositions, and the systems they are part of all contribute to the understanding of their formation processes. While these formation processes may be universal, diversity arises from the differences in starting conditions, such as the birth location or time, host star properties, initial disk conditions, and encounters with other planets in the same system. Planet population synthesis modelling can bridge the gap between observations of protoplanetary disks and the fully formed planetary systems seen today (e.g., Mordasini et al., 2009; Benz et al., 2014; Chambers, 2018; Turrini et al., 2021).
Studying the chemical composition of planet forming environments and the bulk chemical abundances of the the atmospheres of gas giants, enables us to potentially link protoplanetery disks to the planets they produce. The sulfur-, phosphorus-, oxygen-, nitrogen-, and carbon-bearing species along with hydrogen (SPONCH), when observed in the atmospheres of giant planets, can provide information on planets’ atmospheric processes and provide valuable insights about possible formation histories (Öberg et al., 2011; Turrini et al., 2021; Walsh, 2025). A similar logic applies to the elemental budget of rocky planets, super-Earths, or giant planet cores.
@@
In the ISM, sulfur exists almost entirely as an atomic gas (Jenkins, 2009), but in protoplanetary disks, it partitions between volatile molecules (e.g. $\mathrm{H_{2}S}$, CS, SO) and predominantly refractory condensates (such as FeS, MgS) (Semenov et al., 2018; Kama et al., 2019). In Figure 1, we show the evolution of sulfur from a nearly ideal volatile in the diffuse interstellar medium to its almost entirely refractory reservoirs in rocky planetesimals. In this paper, we employ a chemical kinetics model to track the processing of sulfur from a volatile gas into a refractory mineral as a function of time and local physical conditions, without invoking an equilibrium condensation assumption. We suggest this is consistent with evidence from protoplanetary disks (Fuente et al., 2010; Dutrey et al., 2011; Kama et al., 2019) and from solar system comets (Le Roy et al., 2015; Altwegg et al., 2016), which reveals that in addition to a large refractory S reservoir, a significant volatile S reservoir may persist in particular in the outer disk where e.g. cometary ices originate.
Current population synthesis models typically assume 100 $\%$ of sulfur resides in solid condensates (Turrini et al., 2021; Pacetti et al., 2022), making atmospheric sulfur abundances direct tracers of accreted planetesimal/pebble (solid) material and planetary metallicity. However, this assumption is inconsistent with routinely detected sulfur bearing species (CS, SO, $\mathrm{H_{2}S}$) observed in protoplanetary disks by ALMA and Northern Extended Millimeter Array (NOEMA) (Phuong et al., 2018; Booth et al., 2023; Keyte et al., 2024). Studies of Herbig Ae/Be systems reveal that as much as $\approx 11\,$% of all S atoms are in volatile form (Kama et al., 2019), while the diffuse ISM results described above indicate that prior to molecular cloud formation almost all sulfur must be in an atomic gas (Jenkins, 2009). While sulfur will ultimately exist as gas in hot planetary atmospheres regardless of its accretion phase, the bulk sulfur abundance of a planet’s core or atmosphere may vary depending on the relative importance of sulfur inherited as gas, ice, or refractories during different stages of core or envelope accretion. This distinction is preserved in the total atmospheric abundance and can possibly reveal formation pathways.
Figure 1: Volatile (blue) and refractory (red) sulfur reservoirs as a percentage of total locally available sulfur, from the initial interstellar mass reservoir (Jenkins, 2009), through protoplanetary disks (Fuente et al., 2010; Dutrey et al., 2011; Kama et al., 2019), and as seen in solar system comets (Le Roy et al., 2015; Calmonte et al., 2016) and asteroids (Wasson and Kallemeyn, 1988). See Kama et al. (2019) for full details.
@@
Hot Jupiters represent ideal laboratories for testing sulfur chemistry models due to their extreme atmospheric conditions and observational accessibility. Their close-in orbits result in atmospheric temperatures exceeding $1000\rm K$, where sulfur-bearing species achieve chemical equilibrium and become detectable through transmission and emission observations (Hobbs et al., 2021). Critically, at these temperatures, sulfur species remain in the gas phase above the photosphere, making atmospheric abundances observed in transmission spectroscopy a reliable proxy for the complete atmospheric and envelope inventory of sulfur (Fu et al., 2024). This absence of sulfur condensation at observable pressures distinguishes hot Jupiters from cooler gas giant planets where sulfur may condense below the photosphere, obscuring bulk composition.
@@
JWST’s infrared capabilities have already begun revealing substantial sulfur in both Hot Jupiter and Neptune atmospheres that cannot be explained by refractory-only models. The detection of sulfur dioxide in WASP-39b’s atmosphere by JWST represents a breakthrough in exoplanet atmospheric chemistry (Alderson et al., 2023; Tsai et al., 2024; Powell et al., 2024). JWST has also detected hydrogen sulfide ($\mathrm{H_{2}S}$) in HD189733b’s atmosphere, providing the first direct measurement of reduced sulfur species in an exoplanet (Fu et al., 2024), and sulfur dioxide in warm Neptune WASP-107b (Dyrek et al., 2024). These discoveries demonstrate that several reservoirs of sulfur in exoplanet atmospheres are detectable, and may provide new constraints on models of planet formation and atmospheric chemistry. The Ariel mission may eventually substantially expand the sample of exoplanets thus chemically characterised (Tinetti et al., 2022).
@@
Initial Embryo Mass $M_{\rm p,0}$ - $10^{-5}M_{\oplus}$
@@
‘Birth’ Semi-Major Axis $a_{0}$ $0.05-r_{\rm out,0}$ -
Planetesimal-to-solid ratio $f_{\rm ptd}$ - $0.1$
Fragmentation velocity $v_{\rm frag}$ 2.5 $ms^{-1}$ -
@@
Initial disk mass $M_{d,0}$ $0.01-0.1M_{\odot}$ $1\times 10^{-1}M_{\odot}$
Disk wind strength* $f_{\rm w}$ 010-1 $0.8$
@@ 2.1 Disk models & initial conditions
The disk module in sponchpop allows for a variety of 1D, or 1+1D protoplanetary disk models to be implemented. In this work, we make use of 1) viscous and irradiated disks from Chambers (2009); and 2) a viscous, irradiated disk under the influence of an magnetohydrodynamical (MHD) disk wind from Chambers (2018). For the case of the MHD wind disk model, we use fiducial parameters consistent with a fast wind and negligible mass loss due to the winds. In both the MHD wind and viscous irradiated disk models, the boundary between the inner disk (heated through viscous motion) and the flared outer disk (heated through stellar irradiation) are divided by a transition radius that moves inwards over time as the disk mass accretion rate declines. All host stars considered in this study are have solar mass, radius and effective temperature. The initial elemental abundances for $\rm Fe$ and $\rm S$ are taken from Asplund et al. (2009).
@@ 2.2.3 Migration
In the early stages of the planet formation process when the protoplanet is not yet, or never will, reach a mass significant enough to perturb the gas disk, it undergoes Type I migration, wherein the inwards drift of a disk-embedded planet is determined by the interaction between the disk and the planet. Inward migration in this case determined by an imbalanced torque from Lindblad resonances within the planet’s orbit. The migration prescription utilised is from Tanaka et al. (2002). This formulation is used as it assumes the protoplanetary disk is locally vertically isothermal, which is the case for our disk models. Using the total net torque (Lindblad and co-rotation) for a 2D disk,
@@ 2.3 Chemistry
Figure 6: Formation tracks of planets in the viscous irradiated disk with an alpha viscosity value of $\alpha=0.7\times 10^{-3}$ and starting mass of $M_{\rm disk,0}=0.075M_{\odot}$. All other parameters are fiducial. The solid, dashed and dotted parts of each track represent the solid accretion, envelope contraction and runaway gas accretion, respectively. The marker colour denotes the final sulfur mass fraction in the envelope (outer ring) and core (solid centre) respectively, and planets with negligible sulfur content are marked in white (values below $10^{-10}$). The blue shaded region represents the minimum and maximum pebble isolation mass. Planet markers that are empty are planets that have no/negligible sulfur in both the core and envelope. This set of planets is generated with a $p_{\rm ratio}=0$, to demonstrate spatial dependence on sulfur abundance without supplementation from refractory sulfur reservoirs.
@@ 3 Results
Figure 9: Violin plots showing the non-zero sulfur mass fractions in the cores and envelopes of giant planets (0.2 – 2  $M_{\rm J}$) across all three disk models. Each row corresponds to one of the three values considered for the ratio of planetesimals ablated in the planets’ atmospheres. Two cases of $p_{\rm ratio}=1$ are featured: one where sulfur is only considered in the refractory carried by planetesimals, and one where the chemical kinetics described in Section 2.3 are included, therefore including both gas-phase and refractory sulfur reservoirs (this is also the model used for $p_{\rm ratio}=0$ and $0.5$)
. Jupiter’s atmospheric sulfur mass fraction is marked in its appropriate mass category for reference.
@@ 4.2 Disk models
Statistical robustness varies across the disk models considered. While the viscous and viscous-irradiated models each produce hundreds of gas giants, the MHD wind model generates a relatively small number of gas giants ($\leq 50$ for each $p_{\rm ratio}$ case). When MHD results are omitted from our analysis, the final distributions remain indistinguishable, indicating that our giant planet populations are dominated by the other two disk evolution models. The observed exoplanet sample may have contributions from different types of disks, where some types have a higher and others a lower formation likelihood of gas giants.
@@
4.3 FeS formation model assumptions
We have assumed an FeS formation rate corresponding to a 1mm grain size. For a fixed mass of dust, this sets the surface area available to convert gas-phase H2S into solid FeS. While the FeS formation timescale is sensitive to the solid Fe mass per unit reactive surface area, the spatial extent of the ‘sulfur desert’ is primarily controlled by temperature-dependent sublimation and freeze-out processes rather than grain size variations. Smaller grains present in the ISM ($\sim\mathrm{0.1\,\mu m}$) would increase reactive surface area and potentially expand the sulfur desert, while larger 1 cm pebbles would have the opposite effect. Compared to the dramatic temperature effects governing sublimation and freeze-out, grain size variation provides less significant changes in FeS spatial distribution.
[…]
The assumption of location and temperature-dependent planetesimal composition initialized at time $t=0$ maximizes the solid ‘sulfur desert’ effect in our models, particularly at higher $p_{\rm ratio}$ values, since planetesimals do not experience the same radial drift as smaller grains and pebbles. This approach reveals the spatial dependence of sulfur partitioning that we aim to investigate. In contrast, assuming a uniform distribution of sulfur-rich planetesimals across the entire disk would mitigate this location-dependent sulfur chemistry effect. Indeed, other planet population synthesis models that treat sulfur as uniformly refractory throughout the disk (Turrini et al., 2021; Pacetti et al., 2022) do not reproduce the sulfur desert phenomenon we identify here.
We have furthermore neglected any higher-order effects in the H2S–FeS gas-grain system, such as deeper Fe-layers being increasingly masked from surface-level sulfur by the growing FeS layer (Lauretta et al., 1996). This would lengthen the timescale for all S to be incorporated in FeS, but due to the exponential temperature dependence of the overall reaction rate as well as the potential differences between the laboratory test samples and real protoplanetary grains, we have chosen to neglect these effects in the current study.
@@ 4.4 Final sulfur abundances
In the $p_{\rm ratio}=0$ case, the maximum sulfur abundance achieved by a gas giant ($\geq 50M_{\oplus}$) was 1.4 times the present day solar abundance. The maximums reached by the $p_{\rm ratio}=0.5$ and $p_{\rm ratio}=1$ were 5.9 and 9.5 times solar, respectively. The most similar planet to Jupiter generated in our models (with a mass of $321.47\rm M_{\oplus}$ and final semi-major axis of $5.16\rm AU$) produced a final envelope sulfur abundance of 1.37 times the solar sulfur abundance. In order to achieve Jupiter’s sulfur abundance of $4.45\times 10^{-5}$, the planet would need an additional $0.0165M_{\oplus}$ of sulfur in its atmosphere, whether through bombardment of sulfur-rich planetesimals, late stage cometary bombardment, or significant core-envelope mixing. However, Jupiter’s enhanced sulfur abundance is able to be met by comparably sized cold gas giant planets in the case of $p_{\rm ratio}=0.5$ if accreting sulfur-rich planetesimals beyond the $\mathrm{H_{2}S}$ iceline. This only true of planets that have migrated into regions of the disk that have not been depleted of sulfur rich planetesimals during the core accretion stage. This is dependant on the alpha viscosity of the disk, which heavily influences the migration speed of the planet, particularly during Type II migration.
@@ 4.5 Solar System object comparison
The extreme sulfur enhancements observed in Uranus ($31\times$ solar) and Neptune ($46\times$ solar) remain beyond the reach of our models, which achieve maximum enrichments of $\sim 9.5\times$ the solar value even with complete planetesimal ablation. These ice giant abundances likely require additional enrichment mechanisms not captured in our 3 Myr formation time-frame. Cometary bombardment during later evolutionary phases represents a promising avenue for achieving such extreme sulfur abundances, as comets are highly sulfur-rich in both volatile $\mathrm{H_{2}S}$ and refractory (Calmonte et al., 2016), and would ablate efficiently in ice giant atmospheres. While our population synthesis focuses on early formation processes around solar analogues rather than specifically reproducing Solar System architecture, future work incorporating extended evolutionary timescales and cometary contributions could address this discrepancy.
@@ 5 Conclusions
    Sulfur content in planet cores and rocky planets is sensitive to a coupled treatment of volatile and refractory sulfur and is intrinsically related to the formation and migration history. Cores born in the inner few AU ($\leq 3\rm AU$ in our fiducial model) lack refractory sulfur, as the temperature range for efficient FeS formation or retention is conducive to a much smaller region of the disk than previously assumed. Treating sulfur as only being carried in refractories neglects the solid sulfur ‘desert’ which our models predict, where neither efficient iron sulfide production nor $\mathrm{H_{2}S}$ freeze-out is relevant. This fundamentally changes sulfur core budget predictions for gas giant cores and terrestrial planets. Refractory sulfur in inner disk cores would need a later FeS formation time from replenished H2S gas or the inward migration of S-rich planetesimals or cores from further out.
@@
    Atmospheric sulfur abundances indicate formation and migration histories for planets with substantial envelopes. Super-solar atmospheric sulfur requires formation beyond the $\mathrm{H_{2}S}$ iceline as well as efficient planetesimal ablation in the outer disk after the onset of runaway gas accretion. Significant sulfur depletion ($10^{-8}$) indicates either formation in volatile sulfur depleted zones due to efficient refractory S conversion, or lack of access to sulfur bearing planetesimals for late-stage infall. This formation diagnostic can compliment standard C/O analysis based approaches to isolate possible migration histories. The dependence of sulfur chemistry on the local thermal and pressure conditions of the disk midplane offers information regarding the formation histories of gas-giants. Migration histories generate core-envelope compositions that indicate possible relative timing of solid and gas accretion phases, as a coupled treatment of volatile and refractory sulfur can provide information regarding the timing of solid and gas accretion through untouched or depleted planetesimal reservoirs.